OpenStudy (anonymous):

fan/medal to first person to explain and help get answer The volume of a pyramid that fits exactly inside a cube is 18 cubic feet. What is the volume of the cube? (5 points) 6 cubic feet 18 cubic feet 54 cubic feet 72 cubic feet

OpenStudy (anonymous):

I know its not 18 but could you give me the formula and the numbers to plug into to get the answer?

OpenStudy (welshfella):

i messed up then volume of pyramid = (1/3)*h * a where a = area of the base and h = height 18 = ah/3 ah = 54 and what is ah in relation to the cube??

OpenStudy (anonymous):

volume?

OpenStudy (anonymous):

ah is in relation to the volume

OpenStudy (welshfella):

yes area * height = volume

OpenStudy (anonymous):

okay so the answer is 54?

OpenStudy (welshfella):

the area of the cube is the same as area of base of pyramid

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thank you for the help. you did a great job explaining!

OpenStudy (welshfella):

yw